one jet travels 30 mi/h faster than another jet. Both jets left Atlanta, traveling in opposite directions, at 9:00 am. At noon they were 2490 apart. Find their rates.
Answer by mananth(16949)

let the rate of one jet be x
other will be x+30
they are moving away from each other
so add up speeds
x+x+30=2x+30 mph
d=2490
t= 3 hours
3(2x+30)=2490
/3
2x+30=2490/3
2x+30=830
2x=800
x=400
rate of the jets 400 & 430 mph